# Energy-efficient Decentralized Learning via Graph Sparsification

### v1

@@ 1 Introduction
Communication efficiency has been a major consideration in designing learning algorithms, as the cost in communicating model updates, e.g., communication time, bandwidth consumption, and energy consumption, dominates the total operation cost in many application scenarios [1]. Existing works on reducing this cost can be broadly classified into (i) model compression for reducing the cost per communication [3] and (ii) hyperparameter optimization for reducing the number of communications until convergence [4]. The two approaches are orthogonal and can be applied jointly.
@@ 1.1 Related Work
Communication cost reduction. One line of works tried to reduce the amount of data per communication through model compression, e.g., [3]. Another line of works reduced the frequency of communications, e.g., [4]. Later works [9, 10] started to combine model compression and infrequent communications. Recently, it was recognized that better tradeoffs can be achieved by activating subsets of links, e.g., via event-based triggers [9, 10] or predetermined mixing matrices [7, 6]. Our work is closest to [7, 6] by also designing the mixing matrix, but we address a different objective of maximum per-node energy consumption.
@@ 2.1 Decentralized Learning Algorithm
where $\bm{W}^{(k)}=(W^{(k)}_{ij})_{i,j=1}^{m}$ is the $m\times m$ mixing matrix in iteration $k$, and $\eta>0$ is the learning rate. To be consistent with the base topology, $W^{(k)}_{ij}\neq 0$ only if $(i,j)\in E$.
@@
[18, Theorem 2] Let $\bm{J}:={1\over m}\bm{1}\bm{1}^{\top}$. Under assumptions (1)–(3), if there exist a constant $p\in(0,1]$ such that the mixing matrices $\{\bm{W}^{(k)}\}_{k=1}^{K}$, each being symmetric with each row/column summing to one [^2] Originally, [18, Theorem 2] had a stronger assumption that each mixing matrix is doubly stochastic, but we have verified that it suffices to have each row/column summing to one., satisfy
$K(p)$ $:=l(F(\overline{\bm{x}}^{(1)})-F_{\inf})$
Remark: The required number of iterations $K(p)$ depends on the mixing matrix only through the parameter $p$: the larger $p$, the smaller $K(p)$. Originally, [18, Theorem 2] only mandates (2) for the product of $\tau$ mixing matrices, but we consider the case of $\tau=1$ for the tractability of mixing matrix design.
@@ 2.4 Optimization Framework
Lower-level optimization: design link weights $\bm{\alpha}$ to maximize the convergence rate (by maximizing $p$) under a given budget $\Delta$ on the maximum cost per node in each iteration, which results in a required number of iterations of $K(p_{\Delta})$.
Upper-level optimization: design $\Delta$ to minimize the total maximum cost per node ${\Delta}\cdot K(p_{\Delta})$.
@@ 3.1 Simplified Objective
The parameter $p$ that minimizes the required number of iterations $K(p)$ is given by
As (4) is not an explicit function of the mixing matrix, we first relate it to an equivalent quantity that is easier to handle. We can relate $p$ to an explicit function of $\bm{W}$ as follows.
Lemma 3.1.
Lemma 3.1 implies that the lower-level optimization should minimize $\rho$. While it is possible to formulate this minimization in terms of the link weights $\bm{\alpha}$, the resulting optimization problem, with a form similar to [6, (18)], will be intractable due to the presence of non-linear matrix inequality constraint. We thus further simplify the objective as follows.
@@ 3.2 Idea on Leveraging Graph Sparsification
We propose to solve the relaxed lower-level optimization (3.1) based on graph (spectral) sparsification. First, we compute the optimal link weight vector $\bm{\alpha}^{\prime}$ without the budget constraint (7b) by solving the following optimization:
@@
Constraint (8b) ensures $\tilde{\rho}=\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha})\bm{B}^{\top}-\bm{J}\|$ at the optimum, i.e., $\bm{\alpha}^{\prime}$ minimizes $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha})\bm{B}^{\top}-\bm{J}\|$. Optimization (3.2) is a semi-definite programming (SDP) problem that can be solved in polynomial time by existing algorithms [20]. The vector $\bm{\alpha}^{\prime}$ establishes a lower bound on the relaxed objective: if $\bm{\alpha}^{*}$ is the optimal randomized solution for (3.1), then ${\rm I\kern-3.00003ptE}\left[\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{*})\bm{B}^{\top}-\bm{J}\|^{2}\right]\geq\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}-\bm{J}\|^{2}$.
Then, we use a graph sparsification algorithm to sparsify the weighted graph with link weights $\bm{\alpha}^{\prime}$ to satisfy the budget constraint. As $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}-\bm{J}\|^{2}$ $=\max\big{(}(1-\lambda_{2}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}))^{2},$ $(1-\lambda_{m}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}))^{2}\big{)}$, and graph sparsification aims at preserving the original eigenvalues $\left(\lambda_{i}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top})\right)_{i=1}^{m}$ [21], the sparsified link weight vector $\bm{\alpha}_{s}$ is expected to achieve an objective value $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}_{s})\bm{B}^{\top}-\bm{J}\|^{2}$ that approximates the optimal for (3.1).
@@ 3.3.1 Ramanujan-Graph-based Design for a Special Case
Ramanujan graphs can be easily constructed by drawing random $d$-regular graphs until satisfying the Ramanujan definition [23]. By the result of [24], for $d\leq m^{1/3}$, we can generate random $d$-regular graphs in polynomial time. Thus, the above method can efficiently construct a deterministic mixing matrix with guaranteed performance in solving the lower-level optimization for a given budget $\Delta$ such that $d\leq m^{1/3}$.
@@ 3.3.2 Greedy Heuristic for General Case
For the general case, ideally we want to sparsify a weighted graph $G^{\prime}$ with link weights $\bm{\alpha}^{\prime}$ such that the sparsified graph with link weights $\bm{\alpha}_{s}$ will approximate the eigenvalues of $G^{\prime}$ while satisfying the constraint $c_{i}(\bm{\alpha}_{s})\leq\Delta$ for each $i\in V$. While this remains an open problem for general graphs, we propose a greedy heuristic based on the intuition that the importance of a link is reflected in its absolute weight. Specifically, we will find the link $(i,j)$ with the minimum absolute weight according to the solution to (3.2) such that the cost for either node $i$ or node $j$ exceeds the budget $\Delta$, set $\alpha_{(i,j)}=0$, and then find the next link by re-solving (3.2) under this additional constraint, until all the nodes satisfy the budget.
@@ 4 Performance Evaluation
We evaluate the proposed solution for the general case based on a real dataset and the topology of a real wireless network. We defer the evaluation in the special case to Appendix 5.2.
@@ 5 Conclusion
Appendix
5.1 Supporting Proofs
Proof of Lemma 3.1.
One direction $p\leq 1-\rho$ was proved in [6], and we will prove that $p\geq 1-\rho$, or equivalently $1-p\leq\rho$. For this, we rely on the following fact (see [7, Lemma 1] [^6] Although [7, Lemma 1] originally assumed $\bm{W}$ to be doubly stochastic, we have verified that having each row/column summing to one is sufficient.): for any matrix $\bm{A}\in\mathbb{R}^{d\times m}$,
@@ 5.1 Supporting Proofs
5.2 Additional Evaluation Results

### v2

@@ 1 Introduction
Communication efficiency has been a major consideration in designing learning algorithms, as the cost in communicating model updates, e.g., communication time, bandwidth consumption, and energy consumption, dominates the total operation cost in many application scenarios [1]. Existing works on reducing this cost can be broadly classified into (i) model compression for reducing the cost per communication [3, 4] and (ii) hyperparameter optimization for reducing the number of communications until convergence [5]. The two approaches are orthogonal and can be applied jointly.
@@ 1.1 Related Work
Communication cost reduction. One line of works tried to reduce the amount of data per communication through model compression, e.g., [3, 10, 11, 12]. Another line of works reduced the frequency of communications, e.g., [13, 14, 5]. by designing an optimized frequency [13, 14] or an adaptive frequency [5]. A unified analysis of the cost-convergence tradeoff of such solutions was provided in [15]. Later works [16, 17] started to combine model compression and infrequent communications. Recently, it was recognized that better tradeoffs can be achieved by activating subsets of links, e.g., via event-based triggers [16, 17] or predetermined mixing matrices [8, 7]. Our work is closest to [8, 7] by also designing the mixing matrix, but we address a different objective of maximum per-node energy consumption.
@@ 2.1 Decentralized Learning Algorithm
where $\bm{W}^{(k)}=(W^{(k)}_{ij})_{i,j=1}^{m}$ is the $m\times m$ mixing matrix in iteration $k$, and $\eta>0$ is the learning rate. To be consistent with the base topology, $W^{(k)}_{ij}\neq 0$ only if $(i,j)\in E$. Finally we let $\overline{\bm{x}}^{(k)}:={1\over m}\sum_{i=1}^{m}\bm{x}^{(k)}_{i}$.
@@
Let $\bm{J}:={1\over m}\bm{1}\bm{1}^{\top}$ and let $\bm{W}$ be a random symmetric matrix such that each row/column in $\bm{W}$ sums to one. Let $\rho:=\|{\rm I\kern-3.00003ptE}[\bm{W}^{\top}\bm{W}]-\bm{J}\|$. Under assumptions (1)–(3), if each mixing matrix $\bm{W}^{(k)}$ is an i.i.d. copy of $\bm{W}$ and $\rho<1$, then D-PSGD can achieve $\frac{1}{K}\sum_{k=1}^{K}\mathbbm{E}[\|\nabla F(\bm{\overline{x}}^{(k)})\|^{2}]\leq\epsilon$ for any given $\epsilon>0$ ($\overline{\bm{x}}^{(k)}:={1\over m}\sum_{i=1}^{m}\bm{x}^{(k)}_{i}$) when the number of iterations reaches
$K(\rho)$ $:=l(F(\overline{\bm{x}}^{(1)})-F_{\inf})$
$\cdot O\left({\hat{\sigma}^{2}\over m\epsilon^{2}}+{\hat{\zeta}\sqrt{M_{1}+1}+\hat{\sigma}\sqrt{1-\rho}\over(1-\rho)\epsilon^{3/2}}+{\sqrt{(M_{2}+1)(M_{1}+1)}\over(1-\rho)\epsilon}\right).$
Remark: The required number of iterations $K(\rho)$ depends on the mixing matrix only through the parameter $\rho$: the smaller $\rho$, the fewer iterations are needed.
The proof of Theorem 2.1 is based on [25, Theorem 2] and included in Appendix A.
@@ 2.4 Optimization Framework
Lower-level optimization: design link weights $\bm{\alpha}$ to maximize the convergence rate (by minimizing $\rho_{\Delta}$) under a given budget $\Delta$ on the maximum cost per node in each iteration, which results in a required number of iterations of $K(\rho_{\Delta})$.
Upper-level optimization: design $\Delta$ to minimize the total maximum cost per node ${\Delta}\cdot K(\rho_{\Delta})$.
@@ 3.1 Simplified Objective
Theorem 2.1 implies that the lower-level optimization should minimize $\rho$. While it is possible to formulate this minimization in terms of the link weights $\bm{\alpha}$, the resulting optimization problem, with a form similar to [7, (18)], will be intractable due to the presence of non-linear matrix inequality constraint. We thus further simplify the objective as follows.
@@ 3.2 Idea on Leveraging Graph Sparsification
We propose to solve the relaxed lower-level optimization (5) based on graph (spectral) sparsification. First, we compute the optimal link weight vector $\bm{\alpha}^{\prime}$ without the budget constraint (5b) by solving the following optimization:
@@
Constraint (6b) ensures $\tilde{\rho}=\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha})\bm{B}^{\top}-\bm{J}\|$ at the optimum, i.e., $\bm{\alpha}^{\prime}$ minimizes $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha})\bm{B}^{\top}-\bm{J}\|$. Optimization (6) is a semi-definite programming (SDP) problem that can be solved in polynomial time by existing algorithms [27]. The vector $\bm{\alpha}^{\prime}$ establishes a lower bound on the relaxed objective: if $\bm{\alpha}^{*}$ is the optimal randomized solution for (5), then ${\rm I\kern-3.00003ptE}\left[\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{*})\bm{B}^{\top}-\bm{J}\|^{2}\right]\geq\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}-\bm{J}\|^{2}$.
Then, we use a graph sparsification algorithm to sparsify the weighted graph with link weights $\bm{\alpha}^{\prime}$ to satisfy the budget constraint. As $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}-\bm{J}\|^{2}$ $=\max\big{(}(1-\lambda_{2}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}))^{2},$ $(1-\lambda_{m}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top}))^{2}\big{)}$, and graph sparsification aims at preserving the original eigenvalues $\left(\lambda_{i}(\bm{B}\operatorname{diag}(\bm{\alpha}^{\prime})\bm{B}^{\top})\right)_{i=1}^{m}$ [28], the sparsified link weight vector $\bm{\alpha}_{s}$ is expected to achieve an objective value $\|\bm{I}-\bm{B}\operatorname{diag}(\bm{\alpha}_{s})\bm{B}^{\top}-\bm{J}\|^{2}$ that approximates the optimal for (5).
@@ 3.3.1 Ramanujan-Graph-based Design for a Special Case
Ramanujan graphs can be easily constructed by drawing random $d$-regular graphs until satisfying the Ramanujan definition [30]. By the result of [31, 32], for $d=o(\sqrt{m})$, we can generate random $d$-regular graphs in polynomial time. Thus, the above method can efficiently construct a deterministic mixing matrix with guaranteed performance in solving the lower-level optimization for a given budget $\Delta$ such that $d=o(\sqrt{m})$.
3.3.2 Intractability for General Case
We will see that finding a feasible graph sparsifier is computationally hard in the general case. To facilitate the discussions, assume for all $i$ and $j$, $c_{i}^{a}=0$ and $c_{ij}^{b}=1$, so the budget constraint (5b) translates to a maximum degree constraint. For a general base topology $G$ and any fixed $\Delta$, it is not clear whether there exists a feasible Laplacian $\bm{L}$ satisfying the maximum degree constraint (recall that for $\bm{L}$ to be feasible, its convergence parameter $\rho$ for $\bm{W}=\bm{I}-\bm{L}$ must be strictly less than $1$). For example, a subgraph of a star graph needs $m-1$ edges incident to the center to remain connected, and if $\Delta<m-1$, then any subgraph with maximum degree at most $\Delta$ is disconnected, which implies $\lambda_{2}(\bm{L})=0$. Hence, through similar computations as in (3)–(4), for a deterministic $\bm{L}$,
$\rho(\bm{W})=\|\bm{I}-\bm{L}-\bm{J}\|^{2}\geq(1-\lambda_{2}(\bm{L}))^{2}=1.$ (7)
Moreover, in general, the task of determining whether there exists a feasible $\bm{L}$ satisfying the maximum degree constraint is NP-hard because deciding the existence of a connected spanning subgraph with maximum degree no more than $\Delta$ is NP-hard. The following theorem provides NP-hardness for a slightly more general problem; its proof is provided in Appendix A.
Theorem 3.2.
Given a graph $G_{cand}=(V,E_{cand})$, and a degree constraint $d_{i}$ for each vertex $v_{i}\in V$, then it is NP-hard to decide the existence of $E\subseteq E_{cand}$ such that $G=(V,E)$ is a connected graph and $\deg_{G}(v_{i})\leq d_{i}$ for all vertices $v_{i}\in V$.
Finding a feasible graph sparsifier under degree constraints is equivalent to finding a connected spanning subgraph under the same constraints, as
$\rho(\bm{W})=\max\left((1-\lambda_{2}(\bm{L}))^{2},(1-\lambda_{m}(\bm{L}))^{2}\right)$
and we can set the link weights to be small enough so that $\lambda_{m}(\bm{L})<2$, under which $\rho(\bm{W})<1$ if and only if $\lambda_{2}(\bm{L})>0$. Therefore, for a general base topology with general costs and a general budget constraint, it is algorithmically intractable to find a feasible graph sparsifier.
@@ 3.3.3 Greedy Heuristic for General Case
For the general case, ideally we want to sparsify a weighted graph $G^{\prime}$ with link weights $\bm{\alpha}^{\prime}$ such that the sparsified graph with link weights $\bm{\alpha}_{s}$ will approximate the eigenvalues of $G^{\prime}$ while satisfying the constraint $c_{i}(\bm{\alpha}_{s})\leq\Delta$ for each $i\in V$. While this remains an NP-hard problem for general graphs, we propose a greedy heuristic based on the intuition that the importance of a link is reflected in its absolute weight. Specifically, we will find the link $(i,j)$ with the minimum absolute weight according to the solution to (6) such that the cost for either node $i$ or node $j$ exceeds the budget $\Delta$, set $\alpha_{(i,j)}=0$, and then find the next link by re-solving (6) under this additional constraint, until either all the nodes satisfy the budget or the graph becomes disconnected; in the latter case, the algorithm reports failure to find a sparsifier under budget $\Delta$.
@@ 4 Performance Evaluation
We evaluate the proposed solution for the general case based on a real dataset and the topology of a real wireless network. We defer the evaluation in the special case to Appendix B.
@@ 5 Conclusion
Appendix A Supporting Proofs
A.1 Proof of Theorem 2.1
We first recall the following result from [25].
Theorem A.1.
[25, Theorem 2] Let $\bm{J}:={1\over m}\bm{1}\bm{1}^{\top}$. Under assumptions (1)–(3), if there exist a constant $p\in(0,1]$ such that the mixing matrices $\{\bm{W}^{(k)}\}_{k=1}^{K}$, each being symmetric with each row/column summing to one [^5] Originally, [25, Theorem 2] had a stronger assumption that each mixing matrix is doubly stochastic, but we have verified that it suffices to have each row/column summing to one., satisfy
$K(p)$ $:=l(F(\overline{\bm{x}}^{(1)})-F_{\inf})$
Remark: Originally, [25, Theorem 2] only mandates (8) for the product of $\tau$ mixing matrices, but we consider the case of $\tau=1$ for the tractability of mixing matrix design.
Since $\bm{W}^{(1)},\dots,\bm{W}^{(K)}$ are i.i.d. copies of a random matrix $\bm{W}$ in our case, we first rewrite (8) as
Yet (9) is not an explicit function of the mixing matrix, so in the next lemma, we relate it to an equivalent quantity that is an explicit function of $\bm{W}$ and thus easier to handle.
Lemma A.2.
Theorem 2.1 follows from Theorem A.1 and Lemma A.2, so it remains to prove Lemma A.2.
Proof of Lemma A.2.
One direction $p\leq 1-\rho$ was proved in [7], and we will prove that $p\geq 1-\rho$, or equivalently $1-p\leq\rho$. For this, we rely on the following fact (see [8, Lemma 1] [^6] Although [8, Lemma 1] originally assumed $\bm{W}$ to be doubly stochastic, we have verified that having each row/column summing to one is sufficient.): for any matrix $\bm{A}\in\mathbb{R}^{d\times m}$,
@@ Appendix A Supporting Proofs
A.3 Proof of Theorem 3.2
Proof of Theorem 3.2.
It suffices to reduce from the Hamiltonian path problem [35] to our problem. Recall the Hamiltonian path problem for a given graph $G=(V,E)$ is a problem of determining the existence of a path $P\subseteq E$ visiting each vertex $v\in V$ exactly once; such path is referred to as Hamiltonian path.
First, we construct an input graph $G_{cand}$ for each instance $G_{0}=(V_{0},E_{0})$ of the Hamiltonian path problem. Let $V=V_{0},E_{cand}=E_{0}$, and $d_{i}=2$ for all $v_{i}\in V_{0}$.
Second, we show there exists $E\subseteq E_{cand}$ such that if $H=(V_{0},E)$ then $deg_{H}(v_{i})\leq d_{i}$ for each vertex $v_{i}\in V_{0}$ and $\lambda_{2}(\bm{L}_{H})>0$ if and only if $G_{0}$ has a Hamiltonian path.
If there exists a Hamiltonian path $P\subseteq E_{0}$, then since the degree of each vertex in $H=(V_{0},P)$ would be less than or equal to 2, $H$ satisfies the constraint $deg_{H}(v_{i})\leq 2$; moreover, $\lambda_{2}(\bm{L}_{H})>0$ since $H$ is connected.
Conversely, suppose $H=(V_{0},E)$ satisfies $E\subseteq E_{cand}$, $deg_{H}(v_{i})\leq d_{i}$ for each vertex $v_{i}\in V_{0}$ and $\lambda_{2}(\bm{L}_{H})>0$. Then apparently $H$ is connected, and a connected graph with degrees less than or equal to 2 can only be a path or a cycle. A path must be a Hamiltonian path in $G_{0}$ since it connects all the vertices; similarly, a cycle must contain a Hamiltonian path.
The proof is now complete as the Hamiltonian path problem is NP-complete [35]. ∎
Appendix B Additional Evaluation Results